# Tailoring the Hyperparameters of a Wide-Kernel Convolutional Neural Network to Fit Different Bearing Fault Vibration Datasets

### v1

@@ 6 Conclusion
In our introduction, we argued that while there are many successful neural network architectures for bearing fault detection, achieving good performance in real-world applications depends on how well the hyperparameters are tuned for a specific dataset. Benchmark datasets are useful for comparing algorithms but they might not reflect the real-world conditions where the algorithm will be applied, leading to a situation where an algorithm performs well on a benchmark dataset but poorly on a new dataset. Even small changes to hyperparameters can significantly impact the network’s performance and yet there is often no clear guidance on how to choose the best hyperparameters in order to maximise accuracy on new data. We went as far as to say that explaining how hyperparameters affect accuracy on different datasets is more important than focusing on finding ever-newer neural network architectures and we overall argued for a shift in focus from developing new neural network architectures to explaining how to optimise existing ones for real-world applications through better hyperparameter tuning.
@@
Past work showed that the wide-kernel architecture we focus on in this paper is sensitive to changes in the hyperparameters. Our new explorations of a transformer-based architecture and an LSTM architecture for fault detection showed that different kinds of neural network are greatly impacted by the values of the hyperparameters, with the accuracy dropping dramatically when the hyperparameters are set poorly. This serves to emphasise the value of knowing how to set the hyperparameters soundly.
@@
Shifting our focus back to the wide-kernel CNN architecture, a lot of our research focused on the width of the kernel in the first layer. The kernel size is an important hyperparameter and a ‘wide’ kernel in particular has been pointed out in past literature as especially useful for processing raw sensor data. Making the first layer kernel too small causes a drop to 25% average accuracy on CWRU data from 69% when the kernel is large. We had expected the value of a wide kernel to be related to either the sampling rate of the data or the presence of high-frequency noise and yet our experiments showed that neither the sampling rate nor noise is responsible for wide-kernel CNNs performing better than narrow kernel networks. Instead, we are left speculating about what really causes wide kernels to perform better on CWRU data. One possibility is that the wide kernel provides some sort of generic architectural advantage such as regularisation that speeds up training and makes it easier for the training process to converge to a good solution. However, if the advantage of a wide kernel is in some way generic then it is not clear why the phenomenon fails to generalise to SEU data, where a larger kernel performs worse. There could be data properties driving the choice of kernel size which we are simply unaware of.
After considerable exploration our investigations led to some concrete advice on how to tune the hyperparameters: most of them can be set low in order to save computation, whilst the preferred number of filters in layer 1 and in layers 3-5 seems to be 16 or 32. The kernel size in the first layer is the most sensitive to differences between datasets, and it needs to be tuned carefully. Correctly tuned hyperparameters give the final touch to a wide-kernel CNN that performs well on multiple datasets yet also is small and lightweight, making it easier to train and potentially deploy in embedded systems. If performance is lower than expected when generalising to new data then the kernel size in the first layer and the number of filters in layers 3-5 have the highest feature importance and should be tuned first. If the network is still poor-performing on new data then we also suggest an order for tuning the most important hyperparameters individually (Kernel size layer 1, Filters layer 1 and then Filters layers 3-5). Tuning the hyperparameters one-at-a-time saves a huge amount of computation compared to a full grid search and could be more realistic for practitioners to do.
@@
We of course acknowledge there are limitations to our work. Limitations include the fact that we have to restrict our advice only to people analysing raw vibration signals sampled at a fairly high sampling rate. Our insights do not apply to other data, such as data unrelated to (potentially faulty) rotating machines. In other words, our insights are domain-specific. Another disclaimer that should be attached to our work is that although we tried to incorporate as much data as possible, there are only a few benchmark datasets in existence. Even when using seven datasets, we do not have a statistically robust sample, so our conclusions could be influenced by sampling error. It is important to check the results on new data. The accuracy scores reached by neural networks are strongly influenced by how they were trained and so it is also important to contemplate the details of our training approach. One consideration is that we restricted the number of epochs to 100 when training networks and applied early stopping if a network did not improve in validation loss for more than 10 successive epochs. The training procedure must end at some point and so it is necessary to have rules for when to stop in this manner. However, a possible consequence of our stopping rules is that we do not know if networks would have improved further had they been given more time to train. There is a chance that some networks which seem to perform poorly are actually effective but just relatively slow to train. At the same time, there is no reason to assume this is true and the stopping rules might have had no impact on the results. The limitation is that we cannot be completely certain.
Turning to look back at our research as a whole, we see many smaller analyses combining to give a larger picture of the hyperparameters for one specific wide-kernel CNN architecture and also how to set them. We see the hyperparameters impacting how accurate networks are on seven different benchmark datasets. Some hyperparameters can be kept at the same value when shifting between datasets, but the kernel size of the first layer (the so-called ‘wide kernel’ itself) needs to change in response to dataset properties. Multiple experiments with resampling and filtering sought but were ultimately unable to find the data property driving the correct choice of kernel size. Sitting next to these pieces of analysis, we also included an exploration of how the most important hyperparameters interact and which order to tune them in. Framing the entire analysis is our discussion of the importance of understanding hyperparameters when applying neural networks for bearing fault classification. Our main focus was to explain how to set the hyperparameters of a wide-kernel CNN architecture, and we concisely summarised the guidance we can give in that regard, following an extensive data-driven exploration.
CRediT Authorship Contribution Statement
Dan Hudson: Conceptualization, Investigation, Methodology, Project administration, Resources, Software, Validation, Visualization, Writing – original draft, Writing – review and editing
Jurgen van den Hoogen: Conceptualization, Data curation, Investigation, Methodology, Project administration, Resources, Software, Validation, Visualization, Writing – original draft, Writing – review and editing
Martin Atzmueller: Conceptualization, Methodology, Project administration, Supervision, Validation, Writing – review and editing
Funding
This work was supported by funds of zukunft.niedersachsen, Volkswagen Foundation (project “HybrInt – Hybrid Intelligence through Interpretable Artificial Intelligence in Machine Perception and Interaction”).
Declaration of Competing Interest
The authors declare they have no financial and personal relationships with other people or organizations to disclose that could inappropriately influence or bias their work.
Data Availability
The data used in this research (results from the hyperparameter searches) are available on request.

### v2

@@ V Conclusion
In our introduction, we argued that while there are many successful neural network architectures for bearing fault detection, achieving good performance in real-world applications depends on how well the hyperparameters are tuned for a specific dataset. Benchmark datasets are useful for comparing algorithms but they might not reflect the real-world conditions where the algorithm will be applied, leading to a situation where an algorithm performs well on a benchmark but poorly on a new dataset. Even small changes to hyperparameters can significantly impact the network’s performance and yet there is often no clear guidance on how to choose the best hyperparameters in order to maximise accuracy on new data. Our research focused on explaining how hyperparameters affect accuracy on different datasets instead of searching out ever-newer neural network architectures.
@@
Past work showed that the wide-kernel architecture we focus on in this paper is sensitive to changes in the hyperparameters. Our new explorations of a transformer-based architecture and an LSTM architecture for fault detection showed that different kinds of neural network are greatly impacted by the values of the hyperparameters, with the accuracy dropping dramatically when the hyperparameters are set poorly. This serves to emphasise the value of knowing how to set the hyperparameters soundly. Across wide-kernel the CNN, LSTM and transformer architectures we see that bad hyperparameter choices lead to low accuracy when classifying vibration data. Due to differences between datasets, we also found that no single combination of hyperparameter settings works perfectly and there is a benefit to (re)tuning the hyperparameters when moving to new data. This motivated us to identify multiple defaults for the hyperparameters of the wide-kernel CNN architecture, allowing it to be tuned very efficiently for new data.
@@
Keeping our focus exclusively on the wide-kernel CNN architecture, a lot of our research focused on the width of the kernel in the first layer. The kernel size is an important hyperparameter and in past literature a ‘wide’ kernel has been considered useful for processing raw sensor data. Making the first layer kernel too small causes a drop to 25% average accuracy on CWRU data from 69% when the kernel is large. We had expected the value of a wide kernel to be related to either the sampling rate of the data or the presence of high-frequency noise and yet our experiments showed that neither the sampling rate nor noise is responsible for wide-kernel CNNs performing better than narrow kernel networks. Instead, we are left speculating about what really causes wide kernels to perform better on CWRU data. One possibility is that the wide kernel provides some sort of generic architectural advantage such as regularisation that speeds up training or makes it easier for the training process to converge to a good solution. However, if the advantage of a wide kernel is in some way generic then it is not clear why the phenomenon fails to generalise to SEU data, where a larger kernel performs worse. There could be data properties driving the choice of kernel size which we are simply unaware of.
@@
We of course acknowledge there are limitations to our work. Limitations include the fact that we have to restrict our advice only to people analysing raw vibration signals sampled at a fairly high sampling rate. Our insights do not apply to other data, such as data unrelated to (potentially faulty) rotating machines. In other words, our insights are domain-specific. Another disclaimer that should be attached to our work is that although we tried to incorporate as much data as possible, there are only a few benchmark datasets in existence. Even when using seven datasets, we do not have a statistically robust sample, so our conclusions could be influenced by sampling error. It is important to check the results on new data.
The accuracy scores reached by neural networks are strongly influenced by how they were trained.One consideration is that we restricted the number of epochs to 100 when training networks and applied early stopping if a network did not improve in validation loss for more than 10 successive epochs. The training procedure must end at some point and so it is necessary to have rules for when to stop in this manner. However, a possible consequence of our stopping rules is that we do not know if networks would have improved further had they been given more time to train, although there is also no reason to assume this is true.
Turning to look back at our research as a whole, we see many smaller analyses combining to give a larger picture of the hyperparameters for one specific wide-kernel CNN architecture and also how to set them. We see the hyperparameters impacting how accurate networks are on seven different benchmark datasets. To future users, we present “multiple defaults”. Users can try up to ten hyperparameter combinations in order to quickly identify a high-performing configuration of the wide-kernel CNN architecture.
Appendix A Additional Details on Benchmark Datasets
In this appendix we summarise the benchmark datasets. Note that we previously used and described the datasets in [1] and [2].
A-A CWRU Bearing Dataset
The specific choice for this dataset relates to the frequent usage within the fault detection domain  [35, 34, 31, 61, 62, 63], due to its public availability and its structure, which mirrors real-world industrial applications where data is typically not openly accessible. For training purposes, we split the dataset into just 20% train and 80% test data in order to enhance the complexity of the classification task and reduce computation time. For more detailed information, we refer to the original source [60] and the test rig displayed at https://engineering.case.edu/bearingdatacenter/apparatus-and-procedures, along with
A-B Paderborn Dataset
A-C Gearbox
A-D MFPT Dataset
A-E XJTU
A-F UoC
A-G SEU
Appendix B Additional Details on Network Architectures
Here we present the details of the neural network architectures explored in our work. The wide-kernel CNN architecture was previously described in several papers, including [59], [2] and [1].
B-A Long- Short-Term Memory (LSTM) Network
B-B Transformer Architecture
B-C Wide-Kernel Convolutional Neural Network
Our network design consists of two initial convolutional layers, each employing various combinations of stride, kernel size, and filters. Subsequently, a convolutional unit (refer to Figure 10) is replicated three times, resulting in a model with a total of five convolutional layers. In this framework, all three convolutional units share identical hyperparams, based on the rationale that deeper layers should maintain consistent characteristics for optimum learning as established in prior studies [36, 37, 35]. Each convolutional layer is followed by local average pooling, which halves the size of the convolutional output vector by reducing the length $T$ to $\frac{T}{2}$. Finally, the reduced convolutional output is connected to two dense layers, with the last serving as the classification layer.
Figure 10 shows the overall network architecture with two exemplary time series as input.
[…]